# DE-BERT: Distance-Enhanced Early Exiting for BERT based on Prototypical Networks

### v1

@@ III-B1 Update of Class Prototypes
where is to adjust the strength of the update, denotes the update step, and denote the class prototype and class centroid for the th class of the th layer, respectively.
@@ III-B2 Distance-Aware Regularization
We introduce distance-aware regularization (DAR) to train the prototypical networks and learn class prototype representations. Incorporated with the Center Loss [43, 44] that minimizes the intra-class distances, DAR is formulated as the average distance between each sample representation and the prototype of its class:
where denotes the representation of the th sample derived from the th prototypical network, is the ground-truth label of the th sample, and is the number of samples. Note that the class prototypes are fixed when calculating DAR, i.e., no gradients back-propagating through .
@@ III-B3 Training Objective
where , denotes the cross-entropy loss at the th layer, and is a hyper-parameter used to balance the cross-entropy loss and DAR. Since there is no need to exit early at the last layer, we minimize the cross-entropy loss for the classifier at the last (th) layer:
@@
where is the loss of the th layer. Following PABEE [21] and GPFEE [26], the weight of each layer is proportional to its layer number to balance parameter updates across shallow and deep layers. Neither prototypical networks nor internal classifiers share parameters across layers.
@@ III-C1 Entropy-based Correctness Estimation
The entropy provides local information that reflects the model’s confidence level in its predictions for an individual test sample. Following previous works [20, 26, 23], we use the classic normalized entropy to estimate the correctness of early predictions:
@@ III-C2 Distance-based Correctness Estimation
where denotes the sample representation from the th prototypical network. and denote the prototypes of the classes with the highest and the second highest probability scores at the th layer, respectively, and denotes the cosine distance. Intuitively, if , it indicates a consistency with the early prediction, otherwise, it suggests a contradiction with the early prediction result. Therefore, we define a normalized distance ratio to describe the relative relationship between and , which can effectively reflect the correctness of early predictions:
The value of is between and . A lower value indicates a smaller relative to , suggesting a higher correctness of the early prediction. Since and effectively reflect the correctness of early predictions, we focus on these two distances and ignore the distances between the sample and prototypes of other classes for efficiency.
@@ III-C3 Hybrid Exiting Strategy
Based on the above analysis, we propose a hybrid exiting strategy that integrates entropy and distance ratio. We argue the early predictions are sufficiently correct only when the estimation results of prediction correctness derived from both local and global information are sufficiently high. Therefore, we devise a hybrid exiting indicator to improve the estimation of prediction correctness through information fusion, which calculates the harmonic mean of entropy and distance ratio:
where is the normalized distance ratio, and is a parameter used to balance the two factors. The value is always between 0 and 1. A lower value indicates a higher correctness of early prediction from both global and local perspectives. Therefore, once the value falls below the predefined threshold , the inference process is terminated. The inference process of our DE-BERT is summarized in Algorithm 1. For simplicity in expression, entropy and distance ratio mentioned in this paper refer to normalized values.
@@ IV Experiments
In this section, we evaluate the DE-BERT framework’s performance in accelerating inference on the GLUE benchmark. Additionally, we conduct ablation studies to demonstrate the effectiveness of each component within the framework.
@@ IV-A Tasks and Datasets
We select six classification tasks from the GLUE benchmark [45] for experiments, including SST-2, MRPC, QNLI, RTE, QQP, and MNLI. We exclude the STS-B task since it is a regression task. Besides, following the previous studies in [22, 26, 20, 29], we also exclude the WNLI and CoLA tasks. The dataset statistics for all tasks are listed in Table I.
@@ IV-B Baselines
We compare DE-BERT with four groups of baselines:
Backbone: The original backbone provides a reference for all inference acceleration methods, with the speed-up ratio of . We choose the widely used BERT-base [2] as the backbone model for easy comparison.
Budget Exiting: We directly train a BERT-base with layers and denote it as BERT-L. We take and to obtain compressed models with the speed-up ratios of and , respectively. These two baselines enable the exit of all samples at a fixed layer, establishing a lower bound for dynamic early exiting methods as no techniques are employed.
Knowledge Distillation: We select several classic knowledge distillation methods as references, including DistilBERT [46], PD-BERT [47], BERT-PKD [48], and BERT-of-Theseus [49]. These methods leverage different distillation strategies to compress a 12-layer BERT-base model into a 6-layer version, achieving a speed-up ratio.
Early Exiting: We primarily include two groups of early exiting methods for comparison. The first group of baseline methods encompasses all existing exiting strategies, which serves to validate the effectiveness of our DE-BERT framework. It includes the confidence-based strategy DeeBERT [20], the patience-based strategy PABEE [21], the learnable strategy BERxiT [22], and the hybrid strategy PCEE-BERT [23]. The second group includes orthogonal works relevant to our study, GPFEE [26] and LeeBERT [29], offering a benchmark for the performance gains achieved by our DE-BERT within the broader early exiting community. Notably, different from our method which improves exiting strategies, these works focus on improving the training schemes or architectures of internal classifiers for reliable early exiting, addressing early exiting issues from different perspectives. Hence, comparisons with these orthogonal works may not fully reflect the effectiveness of our method.
For fair comparisons, HashEE [24] is not included since it employs a more fine-grained token-level early exiting method, whereas our method focuses on sentence-level early exiting. CascadeBERT [50] is also excluded from the comparison, as it implements early exiting within multiple cascaded complete models instead of a single model with multiple exits employed by our method.
@@ IV-C1 Speed Measurement
Since the actual runtime is unstable across different runs, following [23, 26], we manually adjust the threshold and calculate the corresponding speed-up ratio by comparing the number of executed layers during forward propagation and the total number of layers:
where is the total number of layers and is the number of samples exiting from the th layer. The rationality of this speed measurement is demonstrated in Appendix C-B.
@@ IV-C2 Training
Our implementation is based on Hugging Face’s Transformers [51]. Both the internal classifiers and the prototypical networks are composed of a single linear layer. We choose Center Loss as DAR if not specified and cosine distance as the distance metric. Following the previous work [21, 23, 26], we perform a grid search over learning rates of {1e-5, 2e-5, 3e-5, 5e-5}, batch sizes of , and values in Equation (5) of (the impact of is explored in Section V-A). We set the value in Equation (1) as . The maximum sequence length is fixed at . We employ a linear decay learning rate scheduler and the AdamW [52] optimizer. We conduct experiments on one single RTX3090 GPU with 24GB.
@@ IV-C3 Inference
Following the previous work [23, 26], we use a batch size of during inference, which simulates a common industry scenario where requests from different users arrive one by one. We select in Equation (11) from {0.667, 1.0, 1.5, 2.0, 3.0} for each task. The impact of is explored in Section V-A. We manually adjust the threshold for each task to achieve the speed-up ratios of and , respectively, and focus on the task performance of different methods under these two speed-up ratios.
@@
| BERxiT | 74.5 (2.06) | 84.3/77.2 (1.99) | 85.5 (1.98) | 70.2/88.5 (2.01) | 64.2 (1.97) | 90.2 (1.98) | 79.1 | |
| PCEE-BERT | 81.0 (1.98) | 86.1/80.2 (2.02) | 87.9 (2.04) | 70.8/89.1 (2.02) | 65.7 (1.99) | 91.0 (2.04) | 81.5 | |
| GPFEE | 83.3 (1.96) | 87.0/81.8 (1.98) | 89.8 (1.97) | 71.2/89.4 (2.18) | 64.5 (2.04) | 92.8 (2.02) | 82.5 | |
| LeeBERT | 83.1 (1.97) | 87.1/- (1.97) | - | - | - | 92.6 (1.97) | - | |
| DE-BERT (ours) | 83.2 (2.04) | 86.6/81.5 (1.98) | 90.0 (2.07) | 71.2/89.4 (2.16) | 65.7 (1.99) | 92.5 (2.02) | 82.6 | |
@@
| BERxiT | 60.9 (3.01) | 83.4/75.4 (3.06) | 80.3 (2.89) | 64.3/86.2 (3.14) | 60.7 (2.87) | 84.5 (2.86) | 73.5 | |
| PCEE-BERT | 77.8 (3.08) | 82.8/74.3 (2.98) | 85.6 (3.06) | 69.6/88.3 (2.96) | 63.1 (2.89) | 88.9 (2.92) | 78.8 | |
| GPFEE | 78.4 (2.99) | 84.5/77.7 (2.87) | 87.3 (2.78) | 70.4/89.2 (3.16) | 63.0 (2.88) | 91.1 (2.97) | 80.1 | |
| DE-BERT (ours) | 79.9 (2.99) | 83.8/76.5 (3.01) | 87.0 (3.02) | 70.6/89.3 (3.18) | 63.6 (2.97) | 91.4 (2.98) | 80.3 |
IV-D Overall Performance Comparison
In Table II, we report the classification performance of each early exiting method on the GLUE test sets under the speed-up ratios of 2.0 and 3.0, respectively. To make a fair comparison with the baseline methods, we carefully adjust the threshold to achieve similar speed-up ratios as the baseline methods, and then further compare their task performance. Overall, our method demonstrates a superior performance-efficiency trade-off in nearly all cases when compared to the baseline methods. Concretely, in the context of similar research that improves the exiting strategies during inference, namely BERT-L, DeeBERT [20], PABEE [21], BERxiT [22], and PCEE-BERT [23], our method consistently outperforms these baseline methods across all speed-up ratios by a clear margin, which verifies the effectiveness of our design. Besides, as the speed-up ratio increases, the performance of baseline methods drops dramatically, while the superiority of our method becomes more significant. For instance, under the 2.0 and 3.0 speed-up ratios, our DE-BERT outperforms the state-of-the-art baseline PCEE-BERT [23] by 0.9 and 1.5 points on average across all tasks, respectively. This further indicates that through incorporating class prototypes, our framework can effectively enhance the reliability of exiting decisions, thereby yielding a better trade-off between model performance and efficiency. Notably, our method introduces minimal additional computational or storage costs (see Section V-C). Moreover, in the context of orthogonal works that improve the training schemes or architectures of internal classifiers, namely GPFEE [26] and LeeBERT [29], our method yields comparable results to these baselines. We believe that exploring the combination of our method and these orthogonal works would be an intriguing direction for further research, potentially further boosting the acceleration performance. Besides, we further examine the generality of our method on RoBERTa [3] in Section V-D. We also analyze the impact of different forms of DAR on the acceleration performance of our DE-BERT framework. Please refer to Section V-G for more details.
@@ IV-D Overall Performance Comparison
| DE-BERT (ours) | 83.2 | 90.0 | 89.4 | 92.5 | 88.8 |
In addition, we compare DE-BERT with several classic knowledge distillation methods in Table III. The experimental results consistently demonstrate the superiority of our method. Knowledge distillation methods achieve model compression by distilling the knowledge from a large teacher model to a more compact student model. However, during inference, the compressed model treats samples of different complexities equally, which leads to redundant computations. Instead, our DE-BERT framework enables reliable early exiting through the incorporation of class prototypes. This allows an adaptive number of layers to be executed for each sample based on its complexity, which effectively avoids redundant computations and accelerates the model’s inference speed.
@@ IV-E1 Effect of Distance Ratio
As mentioned in Section III-A, our DE-BERT leverages the global information (distance ratio) to enhance the entropy-based early exiting. To further investigate the effect of distance ratio in exit decision-making, we conduct ablation studies on a representative subset of GLUE to compare the entropy-based exiting strategy with our hybrid exiting strategy. Note that these exiting strategies are applied to the same model trained as illustrated in Section III-B for each task. Consequently, the experimental results encompass the performance improvements brought by DAR, and our focus is solely on the effect of different exiting strategies employed during inference. Fig. 4 shows the performance-efficiency trade-off curves of different exiting strategies. As we can see, our hybrid exiting strategy consistently outperforms the entropy-based strategy by a clear margin under various speed-up ratios across all tasks, confirming the effectiveness of incorporating the distance ratio into the exit decision-making process. This suggests that the distance-based global information can effectively complement the entropy-based local information, which enhances the reliability of exiting decisions, achieving a better trade-off between model performance and inference efficiency. In addition, it is noticeable that the performance improvements brought by the distance ratio appear to be more significant under high speed-up ratios, which confirms the advantage of our DE-BERT framework in high acceleration scenarios. We subsequently analyze the reasons for this observation through experiments. Please refer to Section V-B for details.

### v2

@@ III-B1 Update of Class Prototypes
where is to adjust the strength of the update, denotes the update step, and denote the class prototype and class centroid for the th class at the th layer, respectively.
@@ III-B2 Distance-Aware Regularization
We introduce distance-aware regularization (DAR) to train the prototypical networks and learn class prototype representations. Incorporated with the Center Loss [centerloss, Hu2022DualDC] that minimizes the intra-class distances, DAR is formulated as the average distance between each sample representation and the prototype of its class:
where is the number of samples, and denotes the representation of the th sample derived from the th prototypical network. is the ground-truth label of the th sample, and represents the class prototype corresponding to class at the th layer. The metric is used to compute the cosine distance between and :
where denotes the inner product, and denotes the norm. Note that the class prototypes are fixed when calculating DAR, i.e., no gradients back-propagating through . Algorithm 1 outlines the computation of DAR for each intermediate layer.
@@ III-B3 Training Objective
where , denotes the cross-entropy loss at the th layer, and the regularization coefficient is used to balance the cross-entropy loss and DAR. Since there is no need to exit early at the last layer, we minimize the cross-entropy loss for the classifier at the last (th) layer:
@@
where is the loss at the th layer. Following PABEE [Pabee] and GPFEE [globalpast], the weight of each layer is proportional to its layer number to balance parameter updates across shallow and deep layers. Neither prototypical networks nor internal classifiers share parameters across layers. The training process of DE3-BERT is outlined in Algorithm 2.
@@ III-C1 Entropy-based Correctness Estimation
The entropy provides local information that reflects the model’s confidence level in its predictions for individual test samples. Following previous works [Deebert, globalpast, PCEE], we use the classic normalized entropy to estimate the correctness of early predictions:
@@ III-C2 Distance-based Correctness Estimation
where denotes the sample representation from the th prototypical network. and denote the prototypes of the classes with the highest and the second highest probability scores at the th layer, respectively, and denotes the cosine distance. Intuitively, if , the distance metric result is consistent with the early prediction; otherwise, they are inconsistent. Therefore, we define a normalized distance ratio to describe the relative relationship between and , which effectively reflects the correctness of early predictions:
The value of is between and . A lower value indicates a smaller relative to , suggesting higher correctness of the early prediction. Since and effectively reflect the correctness of early predictions, we focus on these two distances and ignore the distances between the sample and prototypes of other classes for efficiency.
@@ III-C3 Hybrid Exiting Strategy
In this paper, we propose a hybrid exiting strategy that integrates entropy and distance ratio. Specifically, entropy and distance ratio can be regarded as two complementary measures of prediction correctness, derived from local and global information, respectively. Our objective is to enhance the estimation of prediction correctness by jointly considering both perspectives, thereby enabling more reliable and stable exiting decisions. We argue that an early prediction can be deemed sufficiently correct only when the measures of prediction correctness from both local and global perspectives are simultaneously high, i.e., when both entropy and distance ratio values are sufficiently low. Notably, the harmonic mean has been widely adopted for integrating heterogeneous yet complementary measures. For instance, in the computation of the F1 score, precision and recall are fused via harmonic averaging to balance different dimensions and prevent the final decision from being excessively influenced by a single metric. Inspired by this principle, we devise a hybrid exiting indicator, , which computes the harmonic mean of entropy and distance ratio:
where is the normalized distance ratio, and is the fusion coefficient used to balance the two factors. The value is always between 0 and 1. A lower value indicates higher correctness of the early prediction from both global and local perspectives. Therefore, once the value falls below the predefined threshold , the inference process is terminated. The inference process of our DE3-BERT is summarized in Algorithm 3. For simplicity in expression, entropy and distance ratio mentioned in this paper refer to normalized values.
@@ IV Experiments
In this section, we evaluate the DE3-BERT framework’s performance in accelerating inference on the GLUE benchmark. Additionally, we conduct ablation studies to demonstrate the effectiveness of each component within the framework. We further validate the generality of our framework on different backbones.
@@ IV-A Tasks and Datasets
We select six classification tasks from the GLUE benchmark [glue] for experiments, including SST-2, MRPC, QNLI, RTE, QQP, and MNLI. We exclude the STS-B task since it is a regression task. Besides, following the previous studies in [Berxit, globalpast, Deebert, zhu2021leebert], we also exclude the WNLI and CoLA tasks. The dataset statistics for all tasks are listed in Table I.
@@ IV-B Baselines
We compare DE3-BERT with four groups of baselines:
Backbone: The original backbone serves as a reference for all inference acceleration methods, with a speed-up ratio of . We mainly choose the widely used BERT-base [bert] as the backbone model for convincing comparison. We also conduct experiments using RoBERTa-base [abs-1907-11692] on a representative subset of GLUE to evaluate the generality of our method across different backbones.
Budget Exiting: We directly train a BERT-base with layers and denote it as BERT-L. We set to 6 and 4 to achieve and speed-up ratios, respectively. These two baselines force all samples to exit at a fixed layer, providing a lower bound for dynamic early exiting methods since no adaptive mechanisms are applied.
Knowledge Distillation: We select several classic knowledge distillation methods as references, including DistilBERT [distilbert], PD-BERT [wellread], BERT-PKD [PKD], and BERT-of-Theseus [theseus]. These methods leverage different distillation strategies to compress a 12-layer BERT-base model into a 6-layer version, achieving a speed-up ratio.
Early Exiting: We primarily include two groups of early exiting methods for comparison. The first group encompasses all existing exiting strategies, which serves to validate the effectiveness of the proposed distance-enhanced hybrid exiting strategy. It includes the confidence-based strategy DeeBERT [Deebert], the patience-based strategy PABEE [Pabee], the learnable strategy BERxiT [Berxit], and the hybrid strategy PCEE-BERT [PCEE]. The second group includes orthogonal works relevant to our study, i.e., GPFEE [globalpast], LeeBERT [zhu2021leebert], and COSEE [COSEE], offering a benchmark for evaluating the performance of our DE3-BERT within the broader early exiting community. Notably, different from our method which improves exiting strategies, these orthogonal works focus on improving the training schemes or architectures of internal classifiers, addressing early exiting issues from different perspectives. Hence, comparisons with these orthogonal works may not fully represent the effectiveness of the proposed distance-enhanced hybrid exiting strategy.
For fair comparisons, HashEE [hashbased] is not included since it employs a more fine-grained token-level early exiting method, whereas our method focuses on sentence-level early exiting. CascadeBERT [li2021cascadebert] is also excluded from the comparison, as it implements early exiting within multiple cascaded complete models instead of a single model with multiple exits employed by our method.
@@ IV-C1 Speed Measurement
Since the actual runtime is unstable across different runs, following [PCEE, globalpast], we manually adjust the threshold and measure the corresponding speed-up ratio based on the number of layers saved during inference:
where is the total number of layers and is the number of samples exiting from the th layer. We confirm the rationale for this speed measurement in Appendix C-B.
@@ IV-C2 Training
Our implementation is based on Hugging Face’s Transformers [huggingface]. Both the internal classifiers and the prototypical networks are composed of a single linear layer. We choose Center Loss as DAR if not specified and cosine distance as the distance metric. Following the previous work [Pabee, PCEE, globalpast], we perform a grid search over learning rates of {1e-5, 2e-5, 3e-5, 5e-5}, batch sizes of , and values (i.e., the regularization coefficient defined in Equation (6)) of . We set the value in Equation (1) as . The maximum sequence length is fixed at . We employ a linear decay learning rate scheduler and the AdamW [LoshchilovH19] optimizer. We conduct experiments on one single RTX3090 GPU with 24GB.
@@ IV-C3 Inference
Following the previous work [PCEE, globalpast], we use a batch size of during inference, which simulates a common industry scenario where requests from different users arrive one by one. We select the fusion coefficient (in Equation (12)) from {0.667, 1.0, 1.5, 2.0, 3.0} for each task. For fair comparisons, we manually adjust the threshold for each task to achieve similar speed-up ratios as the baseline methods, i.e., and , and then compare their task performance under these two speed-up ratios.
@@
| BERxiT‡ | 74.5 (2.06) | 84.3/77.2 (1.99) | 85.5 (1.98) | 70.2/88.5 (2.01) | 64.2 (1.97) | 90.2 (1.98) | 79.1 | |
| PCEE-BERT‡ | 81.0 (1.98) | 86.1/80.2 (2.02) | 87.9 (2.04) | 70.8/89.1 (2.02) | 65.7 (1.99) | 91.0 (2.04) | 81.5 | |
| GPFEE† | 83.3 (1.96) | 87.0/81.8 (1.98) | 89.8 (1.97) | 71.2/89.4 (2.18) | 64.5 (2.04) | 92.8 (2.02) | 82.5 | |
| LeeBERT† | 83.1 (1.97) | 87.1/- (1.97) | - | - | - | 92.6 (1.97) | - | |
| DE3-BERT (ours) | 83.2 (2.04) | 86.6/81.5 (1.98) | 90.0∗ (2.07) | 71.2/89.4 (2.16) | 65.7 (1.99) | 92.5 (2.02) | 82.6∗ | |
| COSEE† | 83.4 (1.92) | 88.0/82.0 (2.70) | 90.2 (2.56) | 71.4/89.4 (2.01) | 68.7 (1.96) | 93.0 (2.14) | 83.5 | |
| DE-COSEE (ours) | 83.5∘ (1.94) | 88.6/82.2∘ (2.68) | 90.3∘ (2.51) | 71.4/89.4 (2.05) | 68.9∘ (1.98) | 93.0 (2.14) | 83.6∘ | |
@@
| BERxiT‡ | 60.9 (3.01) | 83.4/75.4 (3.06) | 80.3 (2.89) | 64.3/86.2 (3.14) | 60.7 (2.87) | 84.5 (2.86) | 73.5 | |
| PCEE-BERT‡ | 77.8 (3.08) | 82.8/74.3 (2.98) | 85.6 (3.06) | 69.6/88.3 (2.96) | 63.1 (2.89) | 88.9 (2.92) | 78.8 | |
| GPFEE† | 78.4 (2.99) | 84.5/77.7 (2.87) | 87.3 (2.78) | 70.4/89.2 (3.16) | 63.0 (2.88) | 91.1 (2.97) | 80.1 | |
| DE3-BERT (ours) | 79.9∗ (2.99) | 83.8/76.5 (3.01) | 87.0 (3.02) | 70.6/89.3∗ (3.18) | 63.6∗ (2.97) | 91.4∗ (2.98) | 80.3∗ | |
| COSEE‡ | 80.0 (2.98) | 84.2/76.9 (3.05) | 87.2 (3.03) | 70.8/89.4 (3.18) | 64.1 (2.98) | 91.7 (2.97) | 80.6 | |
| DE-COSEE (ours) | 80.3∘ (3.02) | 84.9/77.9∘ (3.03) | 87.5∘ (3.01) | 71.2/89.4∘ (3.17) | 64.9∘ (2.99) | 92.0∘ (3.02) | 81.1∘ |
@@
| DE3-BERT (ours) | 83.2∗ | 90.0∗ | 89.4∗ | 92.5∗ | 88.8∗ |
IV-D Overall Performance Comparison
In Table II, we report the classification performance of each early exiting method on the GLUE test sets at speed-up ratios of 2.0 and 3.0, respectively. Regarding similar research that improves exiting strategies during inference, namely BERT-L, DeeBERT [Deebert], PABEE [Pabee], BERxiT [Berxit], and PCEE-BERT [PCEE], our DE3-BERT consistently outperforms these baseline methods across all speed-up ratios by a clear margin, validating the effectiveness of the proposed distance-enhanced hybrid exiting strategy. Specifically, as the speed-up ratio increases, the performance of these baselines drops dramatically, while the superiority of our method becomes more significant. For instance, at speed-up ratios of 2.0 and 3.0, our DE3-BERT outperforms the competitive baseline PCEE-BERT [PCEE] by an average of 1.1 and 1.5 points across all tasks, respectively. This indicates that by incorporating distance-based global information, our framework effectively improves the reliability of exiting decisions, yielding a better performance-efficiency trade-off for PLMs. Additionally, regarding the remaining orthogonal works that enhance the training schemes or architectures of internal classifiers, our DE3-BERT yields comparable results to GPFEE [globalpast] and LeeBERT [zhu2021leebert]. Since the state-of-the-art baseline COSEE [COSEE] utilizes a calibrated sample weighting mechanism in formulating its training objective, while our DE3-BERT does not employ this technique, a direct comparison would be unfair. To address this issue, we develop a variant DE-COSEE (Distance-Enhanced COSEE), which combines COSEE’s training method with our proposed distance-enhanced hybrid exiting strategy to ensure a fair comparison with the baseline COSEE. The experimental results show that DE-COSEE consistently surpasses the original COSEE in nearly all cases, confirming the effectiveness of our proposed distance-enhanced hybrid exiting strategy for early exiting networks trained with different methods. Furthermore, the one-tailed t-tests verify the statistical significance of our method’s improvements over the top-3 baselines PCEE-BERT [PCEE], GPFEE [globalpast], and COSEE [COSEE] in most cases, providing strong evidence for the superiority of our method in accelerating the inference of PLMs. Notably, our method introduces minimal additional computational and storage costs (see Appendix C-G). Besides, we further explore the impact of hyperparameters and different DAR forms (see Appendices C–E and C–J), and investigate the generality of our method on out-of-distribution (OOD) data and Chinese datasets (see Appendices C–C and C–D).
We also compare DE3-BERT with several classic knowledge distillation methods in Table III. The experimental results consistently demonstrate the superiority of our method, and the one-tailed t-tests further confirm the statistical significance of our method’s improvements over the competitive baselines PD-BERT [wellread] and BERT-of-Theseus [theseus] across all tasks. We attribute this to the static inference process of knowledge distillation methods, where samples of varying complexities pass through the entire model indiscriminately. Instead, our method establishes a reliable dynamic inference process based on sample complexity by incorporating class prototypes. This effectively avoids redundant computations and further enhances the performance-efficiency trade-off of PLMs.
@@ IV-E1 Effect of Distance Ratio
As mentioned in Section III-A, our DE3-BERT leverages the global information (distance ratio) to enhance the entropy-based early exiting. To further investigate the effect of distance ratio in exit decision-making, we conduct ablation studies on a representative subset of GLUE to compare the entropy-based exiting strategy with our hybrid exiting strategy. Note that these exiting strategies are applied to the same model trained as illustrated in Section III-B for each task. Consequently, the experimental results encompass the performance improvements brought by DAR, and our focus is solely on the effect of different exiting strategies employed during inference. Fig. 4 shows the performance-efficiency trade-off curves of different exiting strategies. As we can see, our hybrid exiting strategy consistently outperforms the entropy-based strategy by a clear margin under various speed-up ratios across all tasks, confirming the effectiveness of incorporating the distance ratio into the exit decision-making process. This suggests that the distance-based global information effectively complements the entropy-based local information, thus enhancing the reliability of exiting decisions for better performance-efficiency trade-offs. In addition, it is noticeable that the performance improvements brought by the distance ratio appear to be more significant under high speed-up ratios, demonstrating the advantage of DE3-BERT in high acceleration scenarios. We subsequently analyze the reasons for this observation through experiments (see Appendix C-F).